# Price Experimentation and Interference

### v4

@@
In this paper, we examine the biases that arise when firms run A/B tests on continuous parameters to estimate global treatment effects on performance metrics of interest; we particularly focus on price experiments to measure the price impact on quantity demanded, and on profit. In canonical A/B experimental estimators, biases emerge due to interference between market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to the standard revenue-management pricing problem. This setting highlights a fundamental risk innate to A/B pricing experiments: that the canonical estimator for the expected change in profits, counterintuitively, can have the wrong sign in expectation. In other words, following the guidance of canonical estimators may lead firms to move prices (or fees) in the wrong direction, inadvertently decreasing profits. We introduce a novel debiasing technique for these canonical experiments, requiring only that firms equally split units between treatment and control. We apply these results to a two-sided market model, and demonstrate how the “change of sign” regime depends on market factors such as the supply/demand imbalance, and the price markup. We conclude by calibrating our two-sided market model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners. [^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his critical assistance for our Airbnb empirical case study. We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS Revenue Management and Pricing Conferencea, the INFORMS Annual Meeting, the INFORMS Manufacturing and Service Operations Management Conference, and Airbnb for helpful feedback. Orrie Page was supported by a National Science Foundation Graduate Research Fellowship.
@@ 1 Introduction
Taking advantage of our structural characterization of bias, in Section 6, we construct a novel debiasing method for canonical A/B experiments. In particular, we show that if the experimenter equally allocates units between treatment and control, it is possible to fully debias the canonical estimator by taking advantage of symmetry in the structure of bias. Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below). This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
@@
In Section 8, we conclude the paper by calibrating our two-sided market model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. This exercise simplifies the Airbnb marketplace in many dimensions, but we believe it is useful for offering insights into the prevalence of estimators with the wrong sign in real-world markets. In our investigation, we find that such estimators are plausible in the Airbnb context. Notably, all markets in the study yielded estimators with the wrong sign when we simulated LR experiments, which would be a typical approach for testing pricing changes in practice. Although this analysis is based on a simplified market model, the results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners. We conclude in Section 9.
@@ 2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014, Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
@@
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023). Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation. More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). Roemheld and Rao, (2024) uses an empirical study to suggest that false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility of the sign of the treatment effect estimator being incorrect; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper—those where a continuous variable (such as price) is incrementally varied—as marginal policy changes, and develops a framework to study them.
@@ 3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be adjusted in order to improve the performance metric (a form of policy improvement). We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction $1-q$ to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated to each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
@@
We denote the in-experiment performance metrics for control and treatment goods by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$ respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment values are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units are randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. Notably, this notation abstracts away from the specific experimental design (whether CR or LR). We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1.
@@
$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $10<q<$ 1.
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $10<\alpha<$ 1, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that in practice, experimental estimates typically implicitly rely on the assumption of constant returns in order to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
@@
The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3.2. The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
@@
$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .$ (5)
@@
By Assumption 3.2 and application of the chain rule we get:
@@ 4.2 Experiment: Design, Model, and Estimator
We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction $1-q$ to control. We denote by $p_{0}$ the baseline (control) price and by $p_{1}$ the treatment price. We assume that the experimental demands of control and treated subjects are given by the functions $Q_{0}(p_{0},p_{1},q)$ and $Q_{1}(p_{0},p_{1},q)$, respectively. We further define the functions:
@@
We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1. In addition, $p_{0},p_{1}\geq c$. Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways, commonly assumed in pricing settings: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
@@ 5 Biases: Analysis and Change of Sign
We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $10<q<$ 1.
@@
[…]
@@
$-\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
@@
Second, we note that in our case, we have studied a model where the platform changes the price directly. In practice, another relevant platform parameter that affects profits can be a platform-wide fee. Specifically, in Appendix 12 , we present a model where a platform offers a multitude of heterogeneous goods (potentially at different prices), and earns profit by charging a percentage fee on each transaction. This is a common monetization model for online marketplaces. Viewing the fee as the variable of interest for the platform’s profit optimization, we obtain analogous results to those in this section, including the bias characterization, the possibility of a change of sign, and interpretation of the conditions for a change of sign in terms of appropriate derivatives . This extension suggests our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.
@@ 7.1 Preliminaries
[…]
@@
    $-pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$

### v5

@@
In this paper, we examine the biases that arise when firms run A/B tests on continuous parameters to estimate global treatment effects on performance metrics of interest; we particularly focus on price experiments to measure the price impact on quantity demanded, and on profit. In canonical A/B experimental estimators, biases emerge due to interference between market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to the standard revenue-management pricing problem. This setting highlights a fundamental risk innate to A/B pricing experiments: that the canonical estimator for the expected change in profits, counterintuitively, can have the wrong sign in expectation. In other words, following the guidance of canonical estimators may lead firms to move prices (or fees) in the wrong direction, inadvertently decreasing profits. We introduce a novel debiasing technique for these canonical experiments, requiring only that firms equally split units between treatment and control. We apply these results to a two-sided market model, and demonstrate how the “change of sign” regime depends on market factors such as the supply/demand imbalance, and the price markup. We conclude by calibrating our revenue-management pricing model to published empirical estimates from Airbnb marketplaces, demonstrating that estimators with the wrong sign are not a knife-edge issue, and that they may be prevalent enough to be of concern to practitioners. [^1] We would like to thank Wassim Dhaouadi for his valuable input during the beginning stages of the project, and Andrey Fradkin for his significant assistance on our Airbnb empirical case study. We are also grateful to Eric Aldrich, Alex Deng, Thu Le, Ruben Lobel, Ali Rauh, Dan Schmierer, and seminar participants at Boston University, University of Chile, the MIT Conference on Digital Experimentation, the INFORMS RMP Conference, the INFORMS Annual Meeting, the INFORMS MSOM Conference, and Airbnb for helpful feedback. Orrie Page was supported by a National Science Foundation Graduate Research Fellowship.
@@ 1 Introduction
Taking advantage of our structural characterization of bias, in Section 6, we construct a novel debiasing method for canonical A/B experiments. In particular, we show that if the experimenter equally allocates units between treatment and control, it is possible to fully debias the canonical estimator by taking advantage of symmetry in the structure of bias. Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below). This is a contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require more complicated experimental designs and estimators (Holtz et al., 2025, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
@@
In Section 8, we conclude the paper by calibrating our revenue-management pricing model to empirical estimates from Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. This exercise is useful for offering insights into the plausibility of estimators with the wrong sign in real-world markets. In our investigation, we find that such estimators are quite plausible in the Airbnb context. Notably, all markets in the Farronato and Fradkin, (2022) study demonstrated a susceptibility to estimators with the wrong sign in LR experimentation, which would be the typical approach for testing pricing changes in practice. These results suggest that changes of sign are prevalent enough to be consequential and of concern for practitioners. We conclude in Section 9.
@@ 2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014, Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2025, Candogan et al., 2024), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2023, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), rollout experiments Han et al., (2023), and message passing (Shirani and Bayati, 2024).
@@
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), Roemheld and Rao, (2024), Deng and Le, (2023), Simchi-Levi and Wang, (2023), Delarue and Kleanthis, (2025). Some are less related to our work; for example, Simchi-Levi and Wang, (2023) focuses on designing price experiments to mitigate the tail risk of revenue loss from experimentation, or Delarue and Kleanthis, (2025), which proposes a method for reducing interference bias in matching markets. More related to our paper, both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique to optimize supply-side payments via gradient descent in a centralized marketplace; while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). Roemheld and Rao, (2024) uses an empirical study to suggest that false positives may be inflated due to interference in pricing experiments.
More similar to our findings, Deng and Le, (2023) also notes the possibility of the sign of the treatment effect estimator being incorrect; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. On its part, Li et al., (2025) study A/B experiments comparing recommendation algorithms and shows how symbiosis bias—arising from shared training data between treatment and control—can result in flipping the sign of the estimated global treatment effect. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper—those where a continuous variable (such as price) is incrementally varied—as marginal policy changes, and develops a framework to study them.
@@ 3.2 Experiment: Design, Model, and Estimator
Our work focuses on settings where the function $T$ may be unknown to the firm or platform. Therefore, an experiment is conducted to determine the direction in which the decision variable should be adjusted in order to improve the performance metric (a form of policy improvement). We assume that the platform conducts an experiment with a single treatment condition where a fraction $q$ of the units are randomized to treatment and a fraction 1q1-q to control. We carry out our analysis in a formal “mean field” or continuum regime, where a continuum mass of units are allocated to each group; this allows us to focus on the magnitude and sign of bias (see also Johari et al., (2022), Wager and Xu, (2021) for related works using a similar approach).
@@
We denote the in-experiment performance metrics for control and treatment goods by the functions $T_{0}(x_{0},x_{1},q)$ and $T_{1}(x_{0},x_{1},q)$ respectively. The function $T_{0}(x_{0},x_{1},q)$ (resp., $T_{1}(x_{0},x_{1},q)$) corresponds to the performance metric in the control (resp., treatment) group when the control and treatment values are given by $x_{0}$ and $x_{1}$, respectively, and a fraction $q$ of units are randomized into treatment. The nature of these functions will become more clear in the context of the applications described below. Notably, this notation abstracts away from the specific experimental design (whether CR or LR). [^3] At this level of abstraction, even other experimental designs like switchbacks could be analyzed. We make the following assumption that is kept throughout the paper: {assumption} The functions $T_{0}(x_{0},x_{1},q)$, and $T_{1}(x_{0},x_{1},q)$ are differentiable for all $x_{0},x_{1}\in\mathbb{R}_{+}$ and $0<q<1$.
@@
$T_{0}(x,x,q)/(1-q)=T_{1}(x,x,q)/q=T(x)$, for all $x\in\mathbb{R}_{+}$ and $0<q<1$.
The assumption states that if the values of control and treatment decision variables are the same, then both the scaled control and treatment performance metrics correspond to the market-wide metric. The assumption will play a key role in our analysis. To further understand it, consider $T_{1}$ with an extended domain $0<q\leq 1$ (the argument for $T_{0}$ is symmetric). Now, suppose $T_{1}$ is linear in its third argument over the relevant domain, that is, for any $0<q\leq 1$ and $0<\alpha<1$, $T_{1}(x,x,\alpha q)=\alpha T_{1}(x,x,q)$. If we take $q=1$, then $T_{1}(x,x,\alpha)/\alpha=T_{1}(x,x,1)=T(x)$, where the second equation follows directly by definition. Hence, the assumption is implied by constant (i.e., linear) returns of the performance metric in the fraction randomized. Note that in practice, experimental estimates typically implicitly rely on the assumption of constant returns in order to extrapolate estimates for control and treatment groups to the entire market. For the marketplace models and metrics we study later, this assumption will be satisfied.
@@
The first equation is direct from the definition of $\widehat{GTE}(x_{0},q)$, the second follows by Assumption 3.2, and the third by Assumption 3. The subscript $x$ (resp., $y$) denotes the partial derivative with respect to the first (resp., second) argument of the function.
@@
$=T^{\prime}(x_{0})-\left(T_{1y}(x_{0},x_{0},q)/q-T_{0y}(x_{0},x_{0},q)/(1-q)\right)\ .\$(5)
@@
By Assumption 3 and application of the chain rule we get:
@@ 4.2 Experiment: Design, Model, and Estimator
We assume a single treatment condition in which a fraction $q$ of the subjects are randomized to treatment and a fraction 1q1-q to control. We denote by $p_{0}$ the baseline (control) price and by $p_{1}$ the treatment price. We assume that the experimental demands of control and treated subjects are given by the functions $Q_{0}(p_{0},p_{1},q)$ and $Q_{1}(p_{0},p_{1},q)$, respectively. We further define the functions:
@@
We make the following assumption that implies Assumption 3 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$ and $D_{1}(p_{0},p_{1},q)$ are continuously differentiable in their first two arguments $(p_{0},p_{1})$, for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and all $q\in(0,1)$. In addition, prices satisfy $p_{0},p_{1}\geq c$, where $c\geq 0$, and we assume $0<q<1$.
Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways, commonly assumed in pricing settings: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
@@ 5 Biases: Analysis and Change of Sign
We have that $\widehat{GTE}_{D}(p_{0},q)\leq GTE_{D}(p_{0})$ and $\widehat{GTE}_{\pi}(p_{0},q)\leq GTE_{\pi}(p_{0})$, for all $p_{0}\in\mathbb{R}_{+}$ and $0<q<1$.
@@
$\mbox{Condition (b): \ \ }-\frac{1}{e_{p}(p_{0})}\leq\frac{(p_{0}-c)}{p_{0}}\frac{\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)}{D^{\prime}(p_{0})}:=B(p_{0},q)\ .$ (14)
@@
$\frac{1}{\widehat{e_{p}}(p_{0},q)}\leq\frac{(p_{0}-c)}{p_{0}}\ ,$
@@
Second, we note that in our case, we have studied a model where the platform changes the price directly. In practice, another relevant platform parameter that affects profits can be a platform-wide fee. Specifically, in Appendix 12, we present a model where a platform offers a multitude of heterogeneous goods (potentially at different prices), and earns profit by charging a percentage fee on each transaction. This is a common monetization model for online marketplaces. Viewing the fee as the variable of interest for the platform’s profit optimization, we obtain analogous results to those in this section, including the bias characterization, the possibility of a change of sign, and interpretation of the conditions for a change of sign in terms of appropriate derivatives. This extension suggests that our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.
@@ 7.1 Preliminaries
On the demand side, there is a continuum of customers arriving to the system at rate $\lambda$. When a customer arrives, they form a consideration set consisting of all available listings at that point in time. [^7] It is straightforward to add a process in which customers sample listings to consider with the same probability and independently. We omit this probability to simplify notation. Given the consideration set, customers decide whether or not to book a listing according to a multinomial logit choice model with an outside option of value $\varepsilon>0$. [^8] Note that by assumption we will constrain the value of the outside option $\varepsilon$ to be strictly positive. The value of booking a listing is given by the function $v(p)$. Under this specification, the probability that a customer who arrives in state $s$ books a listing is equal to:
@@
    $pv^{\prime}(p)/v(p)$ is strictly increasing in $p$ for $p\in[c,+\infty).$